OpenStudy (anonymous):

Challenge :D What is 4+5, but not 9?

OpenStudy (anonymous):

Solve this :D

OpenStudy (anonymous):

LOL :D

OpenStudy (anonymous):

Hard eh?

OpenStudy (anonymous):

4+5=9 5=9-4 5=5 :D

OpenStudy (anonymous):

LOL :D

OpenStudy (anonymous):

Wrong

OpenStudy (anonymous):

Hard huh

OpenStudy (anonymous):

I know the answer, tell me when you give up

OpenStudy (anonymous):

No I won't LOL :D Give some hint naa?

OpenStudy (anonymous):

Uh If I give you Hints then it will be obvious

OpenStudy (anonymous):

No just give me one please :D

OpenStudy (anonymous):

:D Uh, you can have variables

OpenStudy (anonymous):

LOL okay :D Wait :P

OpenStudy (anonymous):

It would be 1?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

LOL OMG :o

OpenStudy (anonymous):

This is such an easy question, what math are you taking?

OpenStudy (anonymous):

LOL I'm not taking any Maths :/

OpenStudy (anonymous):

Grade are you in? What?

OpenStudy (anonymous):

Wow I'm only in Algebra 2

OpenStudy (anonymous):

Give up?

OpenStudy (anonymous):

LOL okay :D No i will not :P

OpenStudy (anonymous):

Its going to take you a long time

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

zZZZZZZZZZz

OpenStudy (anonymous):

UMM confusing LOL :/

OpenStudy (anonymous):

Give up?

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

Noooooooooooooo :D

OpenStudy (anonymous):

4+5= not 9 is 4+5=2+7

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What? :o

OpenStudy (anonymous):

I didn't give up LOL :D

OpenStudy (anonymous):

4+9=10-1 LOL xD